# A.I. go by many names: towards a sociotechnical definition of artificial intelligence

### v1

@@ Abstract
Defining artificial intelligence (AI) is a persistent challenge, often muddied by technical ambiguity and varying interpretations. Commonly used definitions heavily emphasize technical properties of AI but neglect the social purpose of it. This essay makes the case for a sociotechnical definition of AI, essential for researchers who require clarity in their work. It explores two primary approaches to define AI: the rationalistic, which focuses on AI as systems that think and act rationally, and the humanistic, which frames AI in terms of its ability to emulate human intelligence. By reconciling these approaches and contrasting them with existing socio-political definitions, the essay proposes a balanced, sociotechnical definition.
@@ Introduction
Table 1 contrasts the scientific conceptualizations of AI with workable definitions that have been proposed in the socio-political discourse in recent years. Notice that most of socio-political definitions exhibit a rationalistic understanding of AI. The most popular definition by the OECD is closely related to Mitchell’s notion of machine learning. Importantly, it emphasizes the systemic properties of applied AI in interaction with its users and environment.
The rationalistic consensus unifies the definitions across regions on a more theoretical level. However, it also exemplifies a pragmatic issue of this approach as the definitions diverge on which specific technologies meet the general definition also practically. This is due to the algorithmic diversity starting at the paradigmatic difference between symbolic and statistical AI. The former—sometimes called good-old-fashioned AI (Buchanan,, 2005)—comprises what Domingos, (2015) describes as tribes of AI researchers focused on logic-based (symbolists) and knowledge-based approaches (analogizers). A common application spawning from this stream of research are expert systems that provide recommendations by following stipulated inference rules connected to a curated base of human knowledge. AI tribes have also formed around the paradigm of statistical AI comprising algorithms that are based on machine learning and deep neural networks (connectionists), evolutionary search- and optimization algorithms such as genetic algorithms (evolutionaries), and using probabilistic modeling techniques such as Bayesian networks (Baysesians).
Given this diversity, contemporary definitions provided by different institutions diverge in the provided examples of technologies usually falling under the umbrella term AI (such as machine learning) and clearly remain non-exhaustive. The proposed European AI Act (European Commission,, 2021) proposes the definition of AI to be complemented with a specific (continuously updated) list of technologies while the U.S. AI Bill of Rights (U.S. Office of Science and Technology Policy,, 2022) makes an intuitive but broad distinction between AI as automated systems and passive computing infrastructure. The definition of the UK’s Alan Turing Institute—possibly due to its intellectual heritage—seems to be one of the only common definitions of AI that acknowledges humanistic elements related to applied AI.
@@
Artificial intelligence (AI) refers to digital machine-based systems that process inputs, infer concepts, and generate outputs for explicit or implicit objectives and to varying degrees of autonomy. The expected purpose of fulfilling these objectives is to automate human labor and/or augment human intelligence.
While one may disagree with this exact sociotechnical definition of AI, it is clear that any such definition must address three central aspects: It must specify a broad technical frame of how AI systems function while delineating AI from more traditional information technologies through the inclusion of concept learning. It must stipulate a clear social purpose of AI, placing the technology in the realm of human-centered societies. It must allow for the fluidity of expectations that users have regarding the behavior that AI technology exhibits while serving its purpose.

### v2

@@ Abstract
Defining artificial intelligence (AI) is a persistent challenge, often muddied by technical ambiguity and varying interpretations. Commonly used definitions heavily emphasize technical properties of AI but neglect the human purpose of it. This essay makes a case for a sociotechnical definition of AI, which is essential for researchers who require clarity in their work. It explores two primary approaches to define AI: the rationalistic, which focuses on AI as systems that think and act rationally, and the humanistic, which frames AI in terms of its ability to emulate human intelligence. By reconciling these approaches and contrasting them with landmark definitions, the essay proposes a sociotechnical definition that includes the three central aspects of i) technical functions, ii) human purpose, and iii) dynamic expectations.
@@ Introduction
Table 1 contrasts the scientific conceptualizations of AI with workable definitions that have been proposed in the socio-political discourse in recent years. Notice that most of them exhibit a rationalistic understanding of AI. The most popular definition by the OECD is closely related to Mitchell’s notion of machine learning. Importantly, it emphasizes the systemic properties of applied AI in interaction with its users and environment.
While the rationalistic consensus unifies the definitions across regions on a more theoretical level, it also exemplifies a pragmatic issue of this approach. The definitions diverge on which specific technologies meet the general definition also practically. This is due to the algorithmic diversity starting at the paradigmatic difference between symbolic and statistical AI. The former—sometimes called good-old-fashioned AI (Buchanan,, 2005)—comprises what Domingos, (2015) describes as tribes of AI researchers focused on logic-based (symbolists) and knowledge-based approaches (analogizers). A common application spawning from this stream of research are expert systems that provide recommendations by following stipulated inference rules connected to a curated base of human knowledge. AI tribes have also formed around the paradigm of statistical AI comprising algorithms that are based on machine learning and deep neural networks (connectionists), evolutionary search- and optimization algorithms such as genetic algorithms (evolutionaries), and using probabilistic modeling techniques such as Bayesian networks (Baysesians).
Given this diversity, contemporary definitions diverge in the provided examples of technologies falling under the umbrella term AI (such as machine learning) and clearly remain non-exhaustive. The proposed European AI Act (European Commission,, 2021) proposes the definition of AI to be complemented with a specific (continuously updated) list of technologies while the U.S. AI Bill of Rights (U.S. Office of Science and Technology Policy,, 2022) makes an intuitive but broad distinction between AI as automated systems and passive computing infrastructure. The definition of the UK’s Alan Turing Institute—possibly due to its intellectual heritage—seems to be one of the only common definitions of AI that acknowledges humanistic elements related to applied AI.
@@
Artificial intelligence (AI) refers to digital machine-based systems that process inputs, infer concepts, and generate outputs for explicit or implicit objectives. It can be expected to perceive and influence its environment to varying degrees of autonomy. The purpose of fulfilling these objectives is to automate human labor and/or augment human intelligence.
While one may disagree with this exact sociotechnical definition of AI, it is clear that any such definition must address three central aspects:
  1. 1.
    It must specify a broad technical frame of how AI systems function while delineating AI from more traditional information technologies through the inclusion of concept learning.
  2. 2.
    It must stipulate a clear human purpose of AI, placing the technology in the realm of human-centered societies.
  3. 3.
    It must allow for the fluidity of expectations that users have regarding the autonomy that AI technology exercises while serving its purpose.